# Dynamical regimes of thermally convective emulsions

### v1

@@ III.1 Statistically steady states
III.2 Transient dynamics and heat flux fluctuations
We complement the steady state characterization of Sec. III.1 with insights into the transient dynamics and heat flux fluctuations. At first, to have a reference case to compare with, we simulated the dynamics of a Newtonian homogeneous fluid under thermal convection in the Rayleigh-Bénard setup. In Fig. 6, we report the time evolution of the Nusselt number, $\operatorname{Nu}$, for different values of $\operatorname{Ra}$. As expected [101], for values of $\operatorname{Ra}$ above a critical value, we observe stable convection in a wide range of values of $\operatorname{Ra}$, roughly between $\operatorname{Ra}\approx 2\times 10^{3}$ and $\operatorname{Ra}\approx 10^{6}$. In this range, the transient dynamics of $\operatorname{Nu}$ shows oscillations whose amplitude decays exponentially over time [113]. Above this upper value, oscillations during transient dynamics are not damped anymore but persist with a periodic pattern, a scenario that is symptomatic of the progressive approach to turbulence [114]. We then select some representative pairs $(\phi,\operatorname{Ra})$ and report the corresponding time evolution of $\operatorname{Nu}$ in Fig. 7. In particular, we consider a dilute emulsion showing a breakup-dominated regime ($\phi=0.16$, panel (a)); a semi-dilute emulsion still showing a breakup-dominated regime ($\phi=0.39$, panel (b)); a concentrated emulsion, just below the jamming point ($\phi=0.68$, panel (c)); a jammed emulsion, just above the jamming point ($\phi=0.79$, panel (d)); a highly packed, jammed emulsion ($\phi=0.84$, panel (e)). Regular oscillations, which are peculiar to the Newtonian homogeneous fluid (see Fig. 6), tend to acquire a stochastic component when $\phi>0$, both in the transient and in the statistically steady state. This randomness is rooted in the presence of finite-sized droplets, which collide while transported by the flow. For fixed $\operatorname{Ra}$, fluctuations in $\operatorname{Nu}$ increase at increasing $\phi$: this is particularly evident when $\operatorname{Ra}\approx 4\times 10^{6}$ and we approach conditions triggering the phase inversion (see panels (a)-(c)). This behavior is similar to what has been observed in a recent study on a Taylor-Couette isothermal system [83], where increasing drag fluctuations are observed while approaching the catastrophic phase inversion. Interestingly, signatures of a beat pattern emerge for $\operatorname{Ra}\approx 4\times 10^{6}$ and are more pronounced for the intermediate values of $\phi$. The combination ($\phi=0.79$, $\operatorname{Ra}\approx 4\times 10^{5}$) reported in panel (d) coincides with the regime explored in Ref. [73], with intermittency characterizing the transient dynamics, where long conductive periods with $\operatorname{Nu}\approx 1$ alternate with intermittent heat bursts where $\operatorname{Nu}\gg 1$, overall resulting in a very long transient dynamics (much longer than the time interval shown in Fig. 7). Furthermore, the transient dynamics observed for $\operatorname{Ra}\approx 4\times 10^{5}$ at changing $\phi$ provides information on how the intermittent regime starts from dumped oscillations (panel (a)), which become more irregular at increasing $\phi$ (panels (b)-(c)), before entering the intermittent regime (panels (d)-(e)). Overall, the intermittent transient dynamics emerges in a narrow range of values of $\operatorname{Ra}$ and $\phi$, corresponding to highly packed emulsions just above the transition from conduction to convection. As discussed in Ref. [73], the intermittency prompts a change in the O/W emulsion structure, whereby heat bursts trigger coalescence events of oil droplets and lead to the phase-inverted W/O emulsion undergoing steady convection. It is noteworthy to observe (see panels (c)-(e) and also Fig. 2) that the coalescence-dominated regime emerges at values of $\operatorname{Ra}$ comparable to those for which the transient intermittency is observed, but at lower values of $\phi$. Following the phase diagram at constant $\operatorname{Ra}\approx 4\times 10^{5}$, starting from high values of $\phi$, the intermittent dynamics is progressively lost as $\phi$ decreases, yet coalescence events are likely to occur. However, since the droplets are less tightly packed, their mobility is enhanced, and the emulsion is more prone to enter statistically steady convective states in shorter times without inducing phase inversion (see panel (c)). In other words, for $\operatorname{Ra}\approx 4\times 10^{5}$ at increasing $\phi$, the coalescence-dominated regime may be seen as a sort of precursor signature of the phase-inverted scenario.
@@ III.2 Transient dynamics and heat flux fluctuations
Overall, one can conclude that the particular transient path that the dynamical evolution selects for a given volume fraction and buoyancy amplitude depends on the initial condition and, in turn, determines the morphology of the statistically steady state. These results are evident, especially in concentrated regimes ($\phi>0.5$) undergoing phase inversion. Such complex phase-inverted states display distinctive morphological and dynamical characteristics compared to an emulsion prepared at the “symmetric” volume fraction $1-\phi$. This “asymmetry” is evident when looking at the transient dynamics at $\operatorname{Ra}\approx 4\times 10^{5}$ for $\phi=0.16$ and $\phi=0.84$ (see panels (a) and (e)), where the dilute emulsion exhibits an almost regular transient state. In contrast, the emulsion with $\phi=0.84$ exhibits the intermittency described in [73]. Nevertheless, there are signatures that the asymmetry is progressively lost as $\operatorname{Ra}$ increases: looking at the data for $\operatorname{Ra}\approx 4\times 10^{6}$, we observe, at least qualitatively, that the behaviors of $\operatorname{Nu}$ as a function of time are much more similar than the behaviors at lower values of $\operatorname{Ra}$, a fact that we will discuss in more detail when studying droplet size statistics in Sec. IV.
@@ IV Droplet size statistics
The phase diagram shown in Fig. 2 reveals that, for sufficiently high values of $\operatorname{Ra}$, only two statistically steady states of convection are observed: a breakup-dominated regime for low-to-moderate values of $\phi$, and a phase inversion regime for higher values of $\phi$. For a given $\phi$, the system settles into a convective state corresponding to one of these two regimes. However, merely analyzing the heat flux properties in these scenarios does not fully capture the structural changes occurring within the emulsion as $\operatorname{Ra}$ varies. Hence, we performed a detailed investigation by examining the droplet size distribution in three representative cases: a semi-dilute emulsion in the breakup-dominated regime ($\phi=0.39$), a yield-stress emulsion undergoing phase inversion ($\phi=0.84$) and a dilute emulsion ($\phi=0.16$), which represents the symmetrical counterpart of the phase-inverted emulsion with $\phi=0.84$. We analyzed the probability distribution functions (PDFs) for the radius, $r$, of the droplets: for this statistical analysis, we consider the oil droplets for the O/W emulsion with $\phi=0.16,0.39$ and the water droplets for the W/O phase-inverted emulsion with $\phi=0.84$. In Fig. 8 we report the PDFs for $\phi=0.39$ and $\phi=0.84$. We also report density maps of the emulsions as $\operatorname{Ra}$ decreases. For $\phi=0.39$, droplets of the oil phase are homogeneously distributed in space, and the breakup events become more frequent as $\operatorname{Ra}$ increases, resulting in a distribution of, mostly, many smaller droplets. For the phase-inverted emulsion, we observe that the phase inversion becomes progressively more and more homogeneous as $\operatorname{Ra}$ increases, with a reduction in the size of the regions occupied by the jammed O/W emulsion. Moreover, the size of the newborn droplets of the water phase changes, producing droplets of smaller size at increasing $\operatorname{Ra}$. Overall, both PDFs share the qualitative feature of a decreasing average value of the droplet size, $r_{m}$, at increasing $\operatorname{Ra}$. This behavior can be understood from a force balance perspective, whereby the turbulent stress imparted by the convective flow to a droplet competes with the surface tension resisting breakup. Put in equations, this means that the velocity fluctuation (squared) at the droplet scale, $(\delta_{r_{m}}u)^{2}$, must balance the Laplace pressure, $\Sigma/r_{m}$, i.e.
@@
Since the Reynolds number is not exceptionally high to justify a fully developed turbulent cascade, the velocity fluctuation can be estimated from the large-scale shear as $\delta_{r_{m}}u\sim(U_{FF}/H)r_{m}$, where $U_{FF}=H/t_{FF}\sim\sqrt{\alpha g\Delta TH}$ is the free-fall velocity [100]. Inserting such expression in (15) we get $(\alpha g\Delta T/H)r_{m}^{3}\sim\Sigma$, whence
@@
This prediction is borne out by analyzing the data on the average droplet size extracted from the PDFs. Results are reported in Fig. 9, where the data for the dilute emulsion at $\phi=0.16$ are also shown. For all three cases, the scaling law in Eq. (16) is verified for almost a decade in $\operatorname{Ra}$. We remark that, in the Hinze’s framework [105], the prefactor in Eq. (15) is related to the critical Weber number for the breakup, $\operatorname{We}_{\text{crit}}$. Such critical value generally depends on the flow pattern around the breaking drops. The intrinsically inhomogeneous and anisotropic nature of flows in the Rayleigh-Bénard setup and the different paths followed to reach the steady state by the emulsions at $\phi=0.16,0.39$ and by the emulsion at $\phi=0.84$ may then lead to different values of $\operatorname{We}_{\text{crit}}$ and account for the different prefactors observed in Fig. 9. In Fig. 10, we show a log-log plot of the PDFs for the same volume fractions considered in Fig. 9. For the emulsion with $\phi=0.39$, we observe a power-law behavior $\sim r^{-10/3}$ for high values of $\operatorname{Ra}$. This agrees with dimensional arguments [115] and with previous studies of drop-laden turbulence [115, 116, 23, 117, 118] and multi-component flows in a Rayleigh-Bénard convection [24]. The same scaling law, however, is not observed for the more dilute emulsion with $\phi=0.16$. To understand why this happens, one must recall that the argument justifying the $-10/3$ scaling, originally devised for the bubbles size distribution in the shallow waters of the ocean, assumes a constant injection of relatively large bubbles being acted upon by turbulent fluctuations, thus breaking up into smaller bubbles which give rise to the PDF tail [119]. To sustain this breakup-driven process, some mechanism that refills the distribution at large sizes is required. In the case of emulsions, where a net influx entraining the dispersed phase from the environment is absent, the supply of large droplets is provided by coalescence events of smaller ones. The system eventually reaches a dynamical equilibrium with a droplet size distribution exhibiting the tail with the $-10/3$ power law. In emulsions, though, because of the effect of the emulsifier, the dispersed phase must not be too dilute to get a significant coalescence rate (which grows with $\phi$) such as to appreciate the emergence of the $-10/3$ scaling. Notice that, this is at odds with unstabilized immiscible liquid mixtures, where the coalescence rate is comparatively high, even at low volume fractions [118]. Finally, echoing the discussion done for Fig. 7, we note that for the cases $\phi=0.16$ and $\phi=0.84$, the PDF tails become progressively similar at increasing $\operatorname{Ra}$, whereas at lower values of $\operatorname{Ra}$ they differentiate.
@@ V Concluding Remarks
We emphasize that all emulsions studied in this work are prepared in a stable state, with a given surface tension at the interface of the droplets and a non-negligible disjoining pressure, effectively preventing droplet coalescence. Hence, the presented scenarios of dynamical regimes under thermal convection have to be considered valid for the employed intensity of disjoining pressure and surface tension: we expect that the boundaries of each region in the phase diagram $\operatorname{Ra}$ vs. $\phi$ will be modified by a variation of these intensities. Actually, the stability induced by the disjoining pressure is a crucial distinction between a simple mixture of two immiscible fluids and a proper emulsion: we conducted preliminary numerical simulations by removing the disjoining pressure in our simulations while leaving the same surface tension at the droplet interface: for very dilute cases ($\phi\rightarrow 0$) the disjoining pressure has no effect at all, as expected, because of the reduced number of droplet interactions; for semi-dilute cases ($\phi\approx 0.4,0.5$) we preliminarily observed that the system structure, as well as the statistical features on droplet size, are manifestly affected. This surely deserves a dedicated study in the future.
We also remark that a peculiar aspect of our study is a thorough characterization of the phase-inverted states; other recent studies focused on the phase inversion process triggered by homogeneous isotropic turbulence with imposed steady forcing [26, 83]. Our analysis in thermally convective flows revealed new properties of the phase inversion close to the thermal instability point, marking the transition from conduction to convection. In particular phase inversion manifests via an irreversible discontinuous jump in the average heat flux $\overline{\operatorname{Nu}}$ as a function of $\operatorname{Ra}$. This fact further supports the view that phase inversion is a complex out-of-equilibrium phenomenon bearing similarity with critical-like systems [83].
From a more general perspective, our simulations offer inspiring insights to envisage novel experiments on thermally convective emulsions and push the field of thermally driven multiphase flows into novel and exciting directions.
@@ Acknowledgements
FP, MS, and MB acknowledge the support of the National Center for HPC, Big Data and Quantum Computing, Project CN_00000013 – CUP E83C22003230001 and CUP B93C22000620006, Mission 4 Component 2 Investment 1.4, funded by the European Union – NextGenerationEU.

### v2

@@ III.2 Transient dynamics and heat flux fluctuations
Furthermore, the transient dynamics observed for $\operatorname{Ra}\approx 4\times 10^{5}$ at changing $\phi$ provides information on how the intermittent regime starts from dumped oscillations (panel (a)), which become more irregular at increasing $\phi$ (panels (b)-(c)), before entering the intermittent regime (panels (d)-(e)). Overall, the intermittent transient dynamics emerges in a narrow range of values of $\operatorname{Ra}$ and $\phi$, corresponding to highly packed emulsions just above the transition from conduction to convection. As discussed in Ref. [73], the intermittency prompts a change in the O/W emulsion structure, whereby heat bursts trigger coalescence events of oil droplets and lead to the phase-inverted W/O emulsion undergoing steady convection. It is noteworthy to observe (see panels (c)-(e) and also Fig. 2) that the coalescence-dominated regime emerges at values of $\operatorname{Ra}$ comparable to those for which the transient intermittency is observed, but at lower values of $\phi$. Following the phase diagram at constant $\operatorname{Ra}\approx 4\times 10^{5}$, starting from high values of $\phi$, the intermittent dynamics is progressively lost as $\phi$ decreases, yet coalescence events are likely to occur. However, since the droplets are less tightly packed, their mobility is enhanced, and the emulsion is more prone to enter statistically steady convective states in shorter times without inducing phase inversion (see panel (c)). In other words, for $\operatorname{Ra}\approx 4\times 10^{5}$ at increasing $\phi$, the coalescence-dominated regime may be seen as a sort of precursor signature of the phase-inverted scenario.
Overall, one can conclude that the particular transient path that the dynamical evolution selects for a given volume fraction and buoyancy amplitude depends on the initial condition and, in turn, determines the morphology of the statistically steady state. These results are evident, especially in concentrated regimes ($\phi>0.5$) undergoing phase inversion. Such complex phase-inverted states display distinctive morphological and dynamical characteristics compared to an emulsion prepared at the “symmetric” volume fraction $1-\phi$. This “asymmetry” is evident when looking at the transient dynamics at $\operatorname{Ra}\approx 4\times 10^{5}$ for $\phi=0.16$ and $\phi=0.84$ (see panels (a) and (e)), where the dilute emulsion exhibits an almost regular transient state. In contrast, the emulsion with $\phi=0.84$ exhibits the intermittency described in [73]. Nevertheless, there are signatures that the asymmetry is progressively lost as $\operatorname{Ra}$ increases: looking at the data for $\operatorname{Ra}\approx 4\times 10^{6}$, we observe, at least qualitatively, that the behaviors of $\operatorname{Nu}$ as a function of time of these two emulsions are much more similar than the behaviors at lower values of $\operatorname{Ra}$, a fact that we will discuss in more detail when studying droplet size statistics in Sec. IV.
@@ IV Droplet size statistics
The phase diagram shown in Fig. 2 reveals that, for sufficiently high values of $\operatorname{Ra}$, only two statistically steady states of convection are observed: a breakup-dominated regime for low-to-moderate values of $\phi$, and a phase inversion regime for higher values of $\phi$. For a given $\phi$, the system settles into a convective state corresponding to one of these two regimes. However, merely analyzing the heat flux properties in these scenarios does not fully capture the structural changes occurring within the emulsion as $\operatorname{Ra}$ varies. Hence, we performed a detailed investigation by examining the droplet size distribution at the statistically steady state in three representative cases: a semi-dilute emulsion in the breakup-dominated regime ($\phi=0.39$), a yield-stress emulsion undergoing phase inversion ($\phi=0.84$) and a dilute emulsion ($\phi=0.16$), which represents the symmetrical counterpart of the phase-inverted emulsion with $\phi=0.84$. We analyzed the probability distribution functions (PDFs) for the radius, $r$, of the droplets: for this statistical analysis, we consider the oil droplets for the O/W emulsion with $\phi=0.16,0.39$ and the water droplets for the W/O phase-inverted emulsion with $\phi=0.84$. In Fig. 8, we report the PDFs for $\phi=0.39$ and $\phi=0.84$. We also report density maps of the emulsions as $\operatorname{Ra}$ decreases. For $\phi=0.39$, droplets of the oil phase are homogeneously distributed in space, and the breakup events become more frequent as $\operatorname{Ra}$ increases, resulting in a distribution of, mostly, many smaller droplets. For the phase-inverted emulsion, we observe that the phase inversion becomes progressively more and more homogeneous as $\operatorname{Ra}$ increases, with a reduction in the size of the regions occupied by the jammed O/W emulsion. Moreover, the size of the newborn droplets of the water phase changes, producing droplets of smaller size at increasing $\operatorname{Ra}$. Overall, both PDFs share the qualitative feature of a decreasing average value of the droplet size, $r_{m}$, at increasing $\operatorname{Ra}$. This behavior can be understood from a force balance perspective, whereby the turbulent stress imparted by the convective flow to a droplet competes with the surface tension resisting breakup. Put in equations, this means that the velocity fluctuation (squared) at the droplet scale, $(\delta_{r_{m}}u)^{2}$, must balance the Laplace pressure, $\Sigma/r_{m}$, i.e.,
@@
Since the Reynolds number is not exceptionally high to justify a fully developed turbulent cascade, the velocity fluctuation can be estimated from the large-scale shear as $\delta_{r_{m}}u\sim(U_{FF}/H)r_{m}$, where $U_{FF}=H/t_{FF}\sim\sqrt{\alpha g\Delta TH}$ is the free-fall velocity [86]. Inserting such expression in Eq. (15), we get $(\alpha g\Delta T/H)r_{m}^{3}\sim\Sigma$, whence
@@
This prediction is borne out by analyzing the data on the average droplet size extracted from the PDFs. Results are reported in Fig. 9, where the data for the dilute emulsion at $\phi=0.16$ are also shown. For all three cases, the scaling law in Eq. (16) is verified for almost a decade in $\operatorname{Ra}$. We remark that, in the Hinze’s framework [110], the prefactor in Eq. (15) is related to the critical Weber number for the breakup, $\operatorname{We}_{\text{crit}}$. Such critical value generally depends on the flow pattern around the breaking drops. The intrinsically inhomogeneous and anisotropic nature of flows in the Rayleigh-Bénard setup and the different paths followed to reach the statistically steady state by the emulsions at $\phi=0.16,0.39$ and by the emulsion at $\phi=0.84$ may then lead to different values of $\operatorname{We}_{\text{crit}}$ and account for the different prefactors observed in Fig. 9. In Fig. 10, we show a log-log plot of the PDFs for the same volume fractions considered in Fig. 9. For the emulsion with $\phi=0.39$, we observe a power-law behavior $\sim r^{-10/3}$ for high values of $\operatorname{Ra}$. This agrees with dimensional arguments [120] and with previous studies of drop-laden turbulence [120, 121, 23, 122, 123] and multi-component flows in a Rayleigh-Bénard convection [24]. The same scaling law, however, is not observed for the more dilute emulsion with $\phi=0.16$. To understand why this happens, one must recall that the argument justifying the $-10/3$ scaling, originally devised for the bubbles size distribution in the superficial layers of the ocean, assumes a constant injection of relatively large bubbles being acted upon by turbulent fluctuations, thus breaking up into smaller bubbles which give rise to the PDF tail [124]. To sustain this breakup-driven process, some mechanism that refills the distribution at large sizes is required. In the case of emulsions, where a net influx entraining the dispersed phase from the environment is absent, the supply of large droplets is provided by coalescence events of smaller ones. The system eventually reaches a dynamical equilibrium with a droplet size distribution exhibiting the tail with the $-10/3$ power law. In emulsions, though, because of the effect of the emulsifier, the dispersed phase must not be too dilute to get a significant coalescence rate (which grows with $\phi$) such as to appreciate the emergence of the $-10/3$ scaling. Notice that, this is at odds with unstabilized immiscible liquid mixtures, where the coalescence rate is comparatively high, even at low volume fractions [123]. Finally, echoing the discussion done for Fig. 7, we note that for the cases $\phi=0.16$ and $\phi=0.84$, the PDF tails become progressively similar at increasing $\operatorname{Ra}$, whereas at lower values of $\operatorname{Ra}$ they differentiate.
@@ V Concluding Remarks
We emphasize that all emulsions studied in this work are prepared in a stable state, with a given surface tension at the interface of the droplets and a non-negligible disjoining pressure, effectively hindering droplet coalescence. As discussed in the text, both the surface tension and the disjoining pressure are considered as independent of the temperature. Hence, the presented scenarios of dynamical regimes under thermal convection have to be considered valid for the employed intensity of disjoining pressure and surface tension: we expect that the boundaries of each region in the phase diagram $\operatorname{Ra}$ vs. $\phi$ will be modified by a variation of these intensities and/or a possible dependence of these quantities on the temperature.
Actually, the stability induced by the disjoining pressure is a crucial distinction between a simple mixture of two immiscible fluids and a proper emulsion: we conducted preliminary numerical simulations by removing the disjoining pressure in our simulations while leaving the same surface tension at the droplet interface: for very dilute cases ($\phi\rightarrow 0$) the disjoining pressure has no effect at all, as expected, because of the reduced number of droplet interactions; for semi-dilute cases ($\phi\approx 0.4,0.5$) we preliminarily observed that the system structure, as well as the statistical features on droplet size, are manifestly affected. This surely deserves a dedicated study in the future. Another interesting aspect to investigate as a future perspective is the possibility to study convective dynamics of emulsions where droplets plasticity at mesoscales is introduced via fluidity-models [68, 125, 126], i.e., effective single fluid models with a non-local constitutive relation between stress and shear rate.
We also remark that a peculiar aspect of our study is a thorough characterization of the phase-inverted states; other recent studies focused on the phase inversion process triggered by homogeneous isotropic turbulence with imposed steady forcing [26, 83]. Our analysis in thermally convective flows revealed new properties of the phase inversion close to the thermal instability point, marking the transition from conduction to convection. In particular, phase inversion manifests via an irreversible discontinuous jump in the average heat flux $\overline{\operatorname{Nu}}$ as a function of $\operatorname{Ra}$. This fact further supports the view that phase inversion is a complex out-of-equilibrium phenomenon bearing similarity with critical-like systems [83].
From a more general perspective, our study offers inspiring ideas for future research. Indeed, we remark that our simulation suite [88] is equipped with a Lagrangian tracking tool for droplets. Hence, one could envisage future research studies where simulations are used to generate ground-truth data for AI studies on the dynamics of deformable objects in complex convective flows. Finally, we argue that our simulations offer inspiring insights to envisage novel experiments on thermally convective emulsions and push the field of thermally driven multiphase flows into novel and exciting directions.
@@ Acknowledgements
FP, MS, and MB acknowledge the support of the National Center for HPC, Big Data and Quantum Computing, Project CN_00000013 – CUP E83C22003230001 and CUP B93C22000620006, Mission 4 Component 2 Investment 1.4, funded by the European Union – NextGenerationEU. This work received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 882340).